# Localize-and-Stitch: Efficient Model Merging via Sparse Task Arithmetic

### v1

@@ 3.2 Localization
        $S_{i}=\operatornamewithlimits{argmin}_{S}\ell_{i}\left(\theta_{\text{pre}}+\sigma(S)\odot\tau_{i}\right)+\lambda\|\sigma(S)\|_{1}$ // make the mask binary $\gamma_{i}=\text{round}(\sigma(S_{i}))$
@@
The objectives are grounded in the findings of [45], which demonstrates that skills developed during finetuning are localizable. Specifically, grafting a small subset of finetuned parameters onto the pretrained model can almost fully recover the performance of the finetuned model. Panigrahi et al. [45] propose the following optimization problem to identify the localized parameters, and we adapt it in the model merging setting. With the constraint on the sparsity level $s$, on the $i$-th task with the loss function $\ell_{i}$ and task vector $\tau_{i}$, we optimize for a binary mask $\gamma_{i}$ such that only adding up the masked portion of the task vector onto the pretrained model performs well on the $i$-th task
$\gamma_{i}=\operatornamewithlimits{argmin}_{\gamma\in\{0,1\}^{d}:\|\gamma\|_{0}\leq s}\ell_{i}(\theta_{\text{pre}}+\gamma\odot\tau_{i}),$
where $\odot$ denotes the element-wise product. For the ease of optimization, we follow [45] to reparametrize the binary mask $\gamma$ as a real-valued vector $S$. To control the sparsity, we additionally relax the $L_{0}$ sparsity constraint to be $L_{1}$. As a result, the optimization is reformulated as
$S_{i}=\operatornamewithlimits{argmin}_{S\in\mathbb{R}^{d}}\ell_{i}\left(\theta_{\text{pre}}+\sigma(S)\odot\tau_{i}\right)+\lambda\|\sigma(S)\|_{1},$ (1)
where $\sigma$ is the sigmoid function, and $\lambda$ controls the strength of the $L_{1}$ regularization. At the end of the optimization, we round $\sigma(S_{i})$ to be binary. A detailed comparison between our localization technique and the one in [45] is presented in Appendix C. Note that the optimization is highly efficient, requiring as few as 8-shot data with 10 epochs of training using SGD. A detailed ablation on the impact of data availability on the mask quality is shown in Section 4.4.
(a) Jaccard similarity of pairwise task masks.
(b) Cosine similarity of masked task vectors.
[…]
Interpretation of task relationships. In Figure 4, we validate that our localization method effectively identifies task-specific regions with minimal overlap. In Figure 4(a), we report the Jaccard similarity of all pairwise task masks, namely for each pair of masks $\gamma_{i}$ and $\gamma_{j}$, we compute $|\gamma_{i}\cap\gamma_{j}|/|\gamma_{i}\cup\gamma_{j}|$. The majority of task pairs exhibit a Jaccard similarity below $5\%$, confirming minimal overlap. For the few pairs with Jaccard similarity larger than $10\%$ (upper left corner of Figure 4(a)), we further compute the cosine similarity of their masked task vectors in Figure 4(b), and find that their cosine similarities are almost 1, indicating high agreement of the parameters within the overlapped regions. Since these four tasks are all sentiment classification tasks, this phenomenon intuitively suggests a shared skill set across the tasks, located in the overlapped regions. We elaborate our resolution for the overlapped regions in Section 3.3
Distribution of localized regions. We analyze the distribution of the localized regions for language tasks in RoBERTa-base models Figure 5, both in terms of the layer index and the transformer components. For the layers, different tasks seem to occupy different layers, although the earlier layers in the network seldomly appear in the localized regions. Interestingly, most of the localized regions concentrate in the LayerNorm parameters. This pattern can possibly be attributed to a covariate shift observed in the finetuning data compared to the pretraining data, necessitating adjustments to the LayerNorm parameters to accommodate this shift. The same plots for GPT2-XL and ViT can be found in Appendix A, and the findings hold true for those models as well.
Localization without validation data. In the rare cases where no labeled data is available, we adopt a similar strategy as the “trim” step in TIES-Merging [67], which selects positions in task vectors with the top-$k\%$ largest magnitudes, i.e., the parameters changed the most during finetuning. We refer to our approach as Dataless Localize-and-Stitch. There are two key differences between Dataless Localize-and-Stitch and TIES: i) We find selecting the top-$5\%$ of parameters is sufficient for our pipeline, compared to the top-$20\%$ recommended by TIES-Merging. ii) We use Stitching described in the next section for merging the localized regions, instead of the “Elect” procedure in TIES. We detail the advantages of those design choices in Appendix B. As shown in Figure 3, to match the performance of localization with validation data, the dataless version typically requires locating larger regions. This expansion is necessary to encapsulate sufficient skills acquired during finetuning, but it also leads to increased task conflicts. Nevertheless, in Section 4, we show that our dataless version still outperforms all other methods that do not require additional validation data.
3.3 Stitching
After obtaining the binary mask for each task, we integrate these masks, and apply them to task vectors to construct the merged model. Given the sparsity, the masks generally activate different positions for different tasks, minimizing overlap. However, in instances where overlaps occur – that is, where multiple tasks share the same mask positions – we address this by averaging the parameters in these regions. Specifically, for each final mask $\gamma_{i}^{\prime}$, the value at the $k$-th position, denoted $\gamma_{i}^{\prime}[k]$, is calculated as the reciprocal of the total number of tasks that have a mask value of 1 at that position; if the original mask value $\gamma_{i}[k]$ is 0, it remains 0 in the processed mask
$\gamma_{i}^{\prime}[k]=\gamma_{i}[k]\bigg{/}\left(\sum_{j=1}^{n}\gamma_{j}[k]\right).$
After we obtain the processed masks $\{\gamma_{i}^{\prime}\}_{i=1}^{n}$, we apply them to the task vectors and stitch the masked task vectors to get the final merged model
@@ 3.3 Stitching
Note that our stitching step does not involve tuning the scaling factors $\alpha$ as other methods mentioned in Section 2, which typically requires grid search or other optimization strategies for tuning. This distinction simplifies our method and avoids the computational overhead. A comparison of runtime is provided in Appendix A.
@@
Remark. In Localize-and-Stitch, the majority of computational overhead lies in the localization step, while the subsequent stitching process is notably efficient. This distribution of workload is ideal because the more intensive localization step is performed separately on each individual finetuned model. This property provides simple extension in continual learning settings: When integrating a new model into the existing merged model (or updating any of the merged models), only the localization step for that new model incurs a cost, followed by stitching. This is in contrast to most model merging methods [32, 30, 67, 68] which necessitate restarting the whole merging process, as the scaling factors of task vectors are tuned or learned based on the performance of the merged model. We provide an experiment of continual learning in Appendix A to illustrate this advantage.
@@ 4 Experiments
We evaluate the performance of Localize-and-Stitch with baselines described in Section 2 under various experimental settings. Our localization step is performed with 64-shot validation data, and the sparsity is chosen to be $1\%$. In the dataless version, the sparsity is chosen to be $5\%$.
@@ 4.1 Merging finetuned encoder-based language models
Following Panigrahi et al. [45], we finetune the RoBERTa-base [38] model on twelve GLUE [59] tasks. Specifically, the dataset suite includes six single-sentence tasks (SST-2 [52], CR [28], MR [44], MPQA [61], TREC [58], SUBJ [43]) and six pairwise-sentence tasks (QNLI [59], SNLI [2], MNLI [62], RTE [59], MRPC [15], QQP [31]). The dataset details can be found in Appendix E.
@@
We present the results in Table 2. The table is structured into three blocks for clarity: the upper block displays the performance of individually finetuned models for each task, the middle block lists algorithms that operate without the need for validation data, whereas the lower block includes algorithms that require validation data. Note that both the middle and lower blocks contain Task arithmetic and TIES because they are applicable with or without data. Both algorithms are able to utilize validation data to tune the merging coefficients $\alpha$, as in $\theta_{\text{merged}}=\theta_{\text{pre}}+\alpha\sum_{i=1}^{n}\tau_{i}$. We follow common practice to search over $\{0.1,0.2,\cdots,1\}$ to obtain the optimal coefficients. When no validation data is available, we use their suggested merging coefficient of 0.4.
@@
From Table 2, regardless of data availability, our approach consistently outperforms other baselines. Notably, the dataless version of our algorithm provides more than $10\%$ performance increase over task arithmetic and surpasses methods that depend on validation data (Fisher merging and AdaMerging), demonstrating its effectiveness. Note that TIES-Merging, although sharing one similar step with our dataless version, performs worse than task arithmetic. This performance decrease is also observed in similar language evaluation settings with similar model size [67]. We identify two possible factors for this phenomenon: i) the larger localized regions of TIES potentially lead to more task conflicts; ii) the sign election mechanism it employs tends to be less effective in overlapping regions that involve only a few tasks, particularly when just two are present. This can lead to suboptimal retention of essential task-specific information. We provide further analysis comparing Dataless Localize-and-Stitch and TIES-Merging in Appendix B.
Table 2: Multi-task performance of merged RoBERTa-base models on twelve NLP tasks.
Task SST-2 CR MR MPQA TREC SUBJ QNLI SNLI MNLI RTE MRPC QQP Average
Simple averaging [64] 0.857 0.851 0.829 0.688 0.304 0.478 0.508 0.452 0.452 0.563 0.311 0.469 0.563
Task artihmetic [30] 0.846 0.856 0.769 0.810 0.156 0.584 0.607 0.538 0.403 0.539 0.822 0.581 0.626
TIES [67] 0.805 0.805 0.728 0.791 0.226 0.549 0.552 0.501 0.379 0.477 0.816 0.572 0.600
Dataless
Localize-and-Stitch
0.909 0.907 0.864 0.821 0.462 0.762 0.558 0.690 0.618 0.688 0.837 0.693 0.734
Task artihmetic [30] 0.885 0.882 0.803 0.829 0.320 0.610 0.620 0.561 0.495 0.656 0.828 0.623 0.675
TIES [67] 0.886 0.88 0.852 0.835 0.226 0.482 0.548 0.359 0.397 0.594 0.794 0.603 0.621
Fisher merging [40] 0.900 0.898 0.837 0.758 0.260 0.546 0.542 0.725 0.652 0.656 0.833 0.677 0.690
RegMean [32] 0.897 0.897 0.847 0.826 0.730 0.791 0.559 0.683 0.568 0.638 0.794 0.642 0.739
AdaMerging [68] 0.850 0.861 0.778 0.815 0.230 0.595 0.612 0.541 0.404 0.547 0.822 0.588 0.637
@@ 4.2 Merging finetuned decoder-based language models
Compared with encoder-only language models, decoder-based language models benefit from increased number of parameters and perform well on complicated generative tasks. We use GPT2-XL [48] as the base model, and obtain three supervised finetuned checkpoints from the Hugging Face model hub [63], each tuned for distinct functionalities: general instruction following, scientific knowledge and truthfulness respectively. Further details about these models are specified in Appendix D.
To assess these models, we use MMLU [26], ARC [9] and TruthfulQA [36] as evaluation datasets for the respective domains. Unlike datasets in the previous section, these are typically used in their entirety for evaluation, without a designated train-test split. However, using these datasets for both evaluation and localization could lead to data leakage. To prevent this, we use data from three surrogate datasets with similar purposes for localization, namely Alpaca [56], GSM8K [10] and HotpotQA [69].
Table 3: Multi-task performance of merged GPT2-XL models on three evaluation benchmarks.
Task MMLU TruthfulQA ARC Average
@@
Simple averaging [64] 0.234 0.390 0.406 0.344
Task arithmetic [30] 0.234 0.390 0.399 0.341
TIES [67] 0.233 0.448 0.310 0.330
Dataless Localize-and-Stitch 0.256 0.394 0.427 0.359
Localize-and-Stitch 0.247 0.388 0.467 0.367
We compare our approach with other methods directly applicable in this setting in Table 3. Both versions of Localize-and-Stitch noticeably outperforms other baselines. The result verifies that even for complex generative tasks, skills can still be localized within tiny regions of finetuned models. Moreover, this shows that good localization performance can be achieved without access to the original finetuning data; using similar data from other sources also suffices. This aligns with the finding from [45] that localized regions exhibit transfer among similar tasks, meaning that a localized region for one task can facilitate performance in related tasks. This further reduces the dependency on data availability, making our approach more versatile. Overall, these findings highlight the capability of Localize-and-Stitch to integrate the strengths from multiple language models, demonstrating its effectiveness across a variety of linguistic challenges.
4.3 Merging finetuned vision models
Following the practice in Ilharco et al. [30], we finetune the CLIP [49] image encoder with the ViT-B/32 [16] architecture on eight image classification tasks, incorporating diverse categories of images such as remote sensing, satellite images and traffic signs. Specifically, the dataset suite includes SUN397 [66], Stanford Cars [33], RESISC45 [6], EuroSAT [25], SVHN [41], GTRSB [53], MNIST [34] and DTD [8]. The details of each dataset can be found in Appendix E.
Table 4: Multi-task performance of merged CLIP ViT-B/32 models on eight vision tasks.
Task SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Average
Simple averaging [64] 0.653 0.634 0.714 0.717 0.642 0.528 0.875 0.501 0.658
Task arithmetic [30] 0.552 0.549 0.667 0.789 0.802 0.697 0.973 0.504 0.692
TIES [67] 0.598 0.586 0.707 0.797 0.862 0.721 0.983 0.542 0.725
Task arithmetic [30] 0.638 0.621 0.720 0.776 0.744 0.651 0.970 0.522 0.701
TIES [67] 0.648 0.629 0.743 0.789 0.831 0.714 0.976 0.562 0.736
Fisher merging [40] 0.686 0.692 0.707 0.664 0.729 0.511 0.879 0.599 0.683
RegMean [32] 0.653 0.635 0.756 0.786 0.781 0.674 0.937 0.520 0.718
AdaMerging [68] 0.645 0.681 0.792 0.938 0.870 0.919 0.975 0.591 0.801
We present the results in Table 4. Similarly, even in the absense of validation data, the dataless version of our approach can outperform methods requiring validation data (Fisher merging and RegMean). When validation data is available, our method also demonstrates competitive performance. Note that AdaMerging, while achieving similar results as ours, imposes more demanding data availability requirement, and incurs higher computational cost. It necessitates entropy minimization across the entire (unlabeled) test set, rendering it approximately 15 times slower than our method.
@@ 4.4 Empirical analysis
Sparsity-performance trade-off. In the localization step, we optimize for two competing objectives: identifying regions containing sufficient finetuned skills, and minimizing the number of parameters involved. Here, we study the trade-off by presenting the performance of our approach on the language tasks at different sparsity levels in Figure 7. Across all models and tasks tested, we observe that a sparsity level around $1\%$ yields the best results using our localization method, whereas dataless localization requires $5-10\%$. When the localized regions are too small to retain adequate finetuned knowledge, the benefit of less overlap is diminished. Conversely, when the localized regions are too large, although the regions possess sufficient finetuned knowledge, the increased overlap among task-specific regions leads to more task interference.
@@
Effect of data availability. We present the performance of our method across various data availability scenarios with a localization region of $1\%$ (Figure 7) on the language tasks. One clear trend is that with more data, the quality of localization improves, resulting in enhanced performance of the merged model. Notably, even with as few as 8-shot data, the merged performance surpasses that of the dataless approach, highlighting its effectiveness under constrained data conditions.
Model compression through localization. Our localization approach enables a compact representation of the finetuned model. In our experiments, we find that localizing only $1\%$ of the total parameters recovers over $99\%$ of the performance achieved by single-task finetuning (full evaluation reported in Appendix A). The efficiency allows us to store only the masked task vectors for each task ($\gamma_{i}\odot\tau_{i}$) instead of the entire finetuned models, without a noticeable loss in performance. Given the sparsity of these masked task vectors, we can store them in Compressed Sparse Row (CSR) format [47, 19], which drastically reduces the model size to about $1\%$ of the original. For example, a RoBERTa-base model, which typically requires $\sim 650$MB of memory to store, can be represented using only a $\sim 7$MB sparse matrix, achieving a memory reduction of $99\%$. Although we still need to store the full pretrained model, this storage overhead will be amortized with more finetuned models. This model compression, combined with the ease of update mentioned in Section 3.3, enables flexible composition of skills from multiple finetuned models with minimal storage and computational overhead.
@@
Avoid forgetting of pretrained knowledge. Pretrained models contain rich and generalizable information, derived from their diverse training data. However, finetuning often incurs catastrophic forgetting of skills in the pretrained model [23, 39], which is carried over when these finetuned models are merged. As our method makes minimal change to the pretrained model, such forgetting is substantially mitigated. For instance, in the vision setting where we localize $1\%$ parameters for each task, the total parameters changed in our merged model compared with the pretrained model is only around $7\%$ for the 8 tasks as a result of minimal overlap. We evaluate the retention of pretrained capabilities with a general vision task that the pretrained CLIP model excels, namely zero-shot ImageNet classification [13], and report the results in Figure 8. Our method most effectively preserves the pretrained performance, while achieves superior performance on the eight merged tasks.

### v2

@@ 3.3 Stitching
        $S_{i}=\operatorname*{arg\,min}_{S}\ell_{i}\left(\theta_{\text{pre}}+\sigma(S)\odot\tau_{i}\right)+\lambda\|\sigma(S)\|_{1}$ // make the mask binary $\gamma_{i}=\text{round}(\sigma(S_{i}))$
@@
The complete algorithm is presented in Algorithm 1. Note that our stitching step does not involve tuning the scaling factors $\alpha$ as other methods mentioned in Section 2, which typically requires grid search or other optimization strategies for tuning. This distinction simplifies our method and avoids the computational overhead. A comparison of runtime is provided in Appendix B.
Remark.
@@
In Localize-and-Stitch, the majority of computational overhead lies in the localization step, while the subsequent stitching process is notably efficient. This distribution of workload is ideal because the more intensive localization step is performed separately on each individual finetuned model. This property provides simple extension in continual learning settings: When integrating a new model into the existing merged model (or updating any of the merged models), only the localization step for that new model incurs a cost, followed by stitching. This is in contrast to most model merging methods Jin et al. (2022); Ilharco et al. (2023); Yadav et al. (2023); Yang et al. (2023) which necessitate restarting the whole merging process, as the scaling factors of task vectors are tuned or learned based on the performance of the merged model. We provide an experiment of continual learning in Section 4.4 to illustrate this advantage.
@@ 4 Experiments
We evaluate Localize-and-Stitch with baselines described in Section 2 under various experimental settings. Our localization step is performed with 64-shot validation data, and the sparsity is chosen to be $1\%$. In the dataless version, the sparsity is chosen to be $5\%$.
Table 2: Multi-task performance of merged RoBERTa-base models on twelve NLP tasks and merged CLIP ViT-B/32 models on eight vision tasks. The reported performance metric is average absolute and normalized (in parentheses) accuracy, with the only exception of MRPC which is evaluated via F1 score due to data imbalance.
Method Validation data Acc/F1 on 12 NLP tasks Acc on 8 vision tasks
Single-task finetuned - 0.811 0.905
Simple averaging (Wortsman et al., 2022a) $0.563_{\;(0.696)}$ $0.658_{\;(0.731)}$
Task artihmetic (Ilharco et al., 2023) $0.626_{\;(0.777)}$ $0.692_{\;(0.758)}$
TIES (Yadav et al., 2023) $0.600_{\;(0.743)}$ $0.725_{\;(0.796)}$
Dataless Localize-and-Stitch $\mathbf{0.734_{\;(0.911)}}$ $\mathbf{0.740_{\;(0.818)}}$
Task arithmetic (Ilharco et al., 2023) $0.675_{\;(0.840)}$ $0.701_{\;(0.778)}$
TIES (Yadav et al., 2023) $0.621_{\;(0.772)}$ $0.736_{\;(0.812)}$
Fisher merging (Matena & Raffel, 2022) $0.690_{\;(0.863)}$ $0.683_{\;(0.763)}$
RegMean (Jin et al., 2022) $0.739_{\;(0.911)}$ $0.718_{\;(0.792)}$
AdaMerging (Yang et al., 2023) $0.637_{\;(0.790)}$ $\mathbf{0.801_{\;(0.880)}}$
Concensus TA (Wang et al., 2024b) $0.715_{\;(0.889)}$ $0.737_{\;(0.810)}$
Consensus TIES (Wang et al., 2024b) $0.695_{\;(0.863)}$ $0.748_{\;(0.822)}$
Localize-and-Stitch $\mathbf{0.759_{\;(0.936)}}$ $\mathbf{0.799_{\;(0.880)}}$
@@ 4.1 Merging finetuned encoder-based language models
Following Panigrahi et al. (2023), we finetune the RoBERTa-base (Liu et al., 2019) model on twelve GLUE (Wang et al., 2018) tasks. Specifically, the dataset suite includes six single-sentence tasks (SST-2 (Socher et al., 2013), CR (Hu & Liu, 2004), MR (Pang & Lee, 2005), MPQA (Wiebe et al., 2005), TREC (Voorhees et al., 1999), SUBJ (Pang & Lee, 2004)) and six pairwise-sentence tasks (QNLI (Wang et al., 2018), SNLI (Bowman et al., 2015), MNLI (Williams et al., 2017), RTE (Wang et al., 2018), MRPC Dolan & Brockett (2005), QQP (Iyer et al., )). The dataset details can be found in Appendix G.
@@
We present the results in Table 2, and leave the detailed per-task results in Table 4 of Appendix A. The table is structured into three blocks for clarity: the upper block displays the performance of individually finetuned models for each task, the middle block lists algorithms that operate without the need for validation data, whereas the lower block includes algorithms that require validation data. Note that both the middle and lower blocks contain Task arithmetic and TIES because they are applicable with or without data. Both algorithms are able to utilize validation data to tune the merging coefficients $\alpha$, as in $\theta_{\text{merged}}=\theta_{\text{pre}}+\alpha\sum_{i=1}^{n}\tau_{i}$. We follow common practice to search over $\{0.1,0.2,\cdots,1\}$ to obtain the optimal coefficients. When no validation data is available, we use their suggested merging coefficient of 0.4.
@@
From Table 2, regardless of data availability, our approach consistently outperforms other baselines. Notably, the dataless version of our algorithm provides more than $10\%$ performance increase over task arithmetic and surpasses methods that depend on validation data (Fisher merging and AdaMerging), demonstrating its effectiveness. Note that TIES-Merging, although sharing one similar step with our dataless version, performs worse than task arithmetic. This performance decrease is also observed in similar language evaluation settings with similar model size Yadav et al. (2023). This phenomenon can be attributed to the two possible factors we identify in Section 3.3: i) the larger localized regions of TIES potentially lead to more task conflicts; ii) the sign election mechanism it employs tends to be less effective in overlapping regions that involve only a few tasks, particularly when just two are present. This can lead to suboptimal retention of essential task-specific information. We provide further analysis comparing Dataless Localize-and-Stitch and TIES-Merging in Appendix C.
4.2 Merging finetuned vision models
Following the practice in Ilharco et al. (2023), we finetune the CLIP (Radford et al., 2021) image encoder with the ViT-B/32 (Dosovitskiy et al., 2021) architecture on eight image classification tasks, incorporating diverse categories of images such as remote sensing, satellite images and traffic signs. Specifically, the dataset suite includes SUN397 (Xiao et al., 2016), Stanford Cars Krause et al. (2013), RESISC45 (Cheng et al., 2017), EuroSAT (Helber et al., 2019), SVHN (Netzer et al., 2011), GTRSB (Stallkamp et al., 2011), MNIST (LeCun et al., 2010) and DTD (Cimpoi et al., 2014). The details of each dataset can be found in Appendix G.
We present the results in Table 2, and leave the detailed per-task results in Table 5 of Appendix A. Similarly, even in the absence of validation data, the dataless version of our approach can outperform methods requiring validation data (Fisher merging and RegMean). When validation data is available, our method also demonstrates competitive performance. Note that AdaMerging, while achieving similar results as ours, imposes more demanding data availability requirement, and incurs higher computational cost. It necessitates entropy minimization across the entire (unlabeled) test set, rendering it approximately 15 times slower than our method.
@@ 4.3 Merging finetuned decoder-based language models
Compared with encoder-only language models, decoder-based language models benefit from increased number of parameters and perform well on complicated generative tasks. We use GPT2-XL (Radford et al., 2019) as the base model, and obtain three supervised finetuned checkpoints from the Hugging Face model hub (Wolf et al., 2020), each tuned for distinct functionalities: general reasoning, scientific knowledge and truthfulness respectively. Further details about these models are specified in Appendix F.
To assess these models, we use MMLU (Hendrycks et al., 2021), ARC (Clark et al., 2018) and TruthfulQA (Lin et al., 2021) as evaluation datasets for the respective domains. Unlike datasets in the previous section, these are typically used in their entirety for evaluation, without a designated train-test split. However, using these datasets for both evaluation and localization could lead to data leakage. To prevent this, we use data from three surrogate datasets with similar purposes for localization, namely Alpaca (Taori et al., 2023), GSM8K (Cobbe et al., 2021) and HotpotQA (Yang et al., 2018).
Table 3: Multi-task performance of merged GPT2-XL models (normalized metrics shown in the parentheses).
Task MMLU (5-shot Acc) TruthfulQA (MC2) ARC (Acc) Average
@@
Simple averaging (Wortsman et al., 2022a) $0.234_{\;(0.857)}$ $0.390_{\;(0.799)}$ $0.406_{\;(0.860)}$ $0.344_{\;(0.839)}$
Task arithmetic (Ilharco et al., 2023) $0.234_{\;(0.857)}$ $0.390_{\;(0.799)}$ $0.399_{\;(0.845)}$ $0.341_{\;(0.834)}$
TIES (Yadav et al., 2023) $0.233_{\;(0.853)}$ $0.448_{\;(0.918)}$ $0.310_{\;(0.657)}$ $0.330_{\;(0.809)}$
Consensus TA (Wang et al., 2024b) $0.234_{\;(0.857)}$ $0.412_{\;(0.844)}$ $0.373_{\;(0.790)}$ $0.340_{\;(0.831)}$
Consensus TIES (Wang et al., 2024b) $0.232_{\;(0.850)}$ $0.395_{\;(0.809)}$ $0.386_{\;(0.818)}$ $0.337_{\;(0.826)}$
Dataless Localize-and-Stitch $0.256_{\;(0.938)}$ $0.394_{\;(0.807)}$ $0.427_{\;(0.905)}$ $0.359_{\;(0.883)}$
Localize-and-Stitch $0.247_{\;(0.905)}$ $0.388_{\;(0.795)}$ $0.467_{\;(0.989)}$ $\mathbf{0.367_{\;(0.896)}}$
We compare our approach with other methods directly applicable in this setting in Table 3. Both versions of Localize-and-Stitch noticeably outperforms other baselines. The result verifies that even for complex generative tasks, skills can still be localized within tiny regions of finetuned models. Moreover, this shows that good localization performance can be achieved without access to the original finetuning data; using similar data from other sources also suffices. This aligns with the finding from (Panigrahi et al., 2023) that localized regions exhibit transfer among similar tasks, meaning that a localized region for one task can facilitate performance in related tasks. This further reduces the dependency on data availability, making our approach more versatile. Overall, these findings highlight the capability of Localize-and-Stitch to integrate the strengths from multiple language models, demonstrating its effectiveness across a variety of linguistic challenges.
@@ 4.4 Empirical analysis
Sparsity-performance trade-off.
In the localization step, we optimize for two competing objectives: identifying regions containing sufficient finetuned skills, and minimizing the number of parameters involved. Here, we study the trade-off by presenting the performance of our approach on the language tasks at different sparsity levels in Figure 7. Across all models and tasks tested, we observe that a sparsity level around $1\%$ yields the best results using our localization method, whereas dataless localization requires $5-10\%$. When the localized regions are too small to retain adequate finetuned knowledge, the benefit of less overlap is diminished. Conversely, when the localized regions are too large, although the regions possess sufficient finetuned knowledge, the increased overlap among task-specific regions leads to more task interference.
@@
Effect of data availability.
We present the performance of our method across various data availability scenarios with a localization region of $1\%$ (Figure 7) on the language tasks. One clear trend is that with more data, the quality of localization improves, resulting in enhanced performance of the merged model. Notably, even with as few as 8-shot data, the merged performance surpasses that of the dataless approach, highlighting its effectiveness under constrained data conditions.
@@
Avoid forgetting of pretrained knowledge.
Pretrained models contain rich and generalizable information, derived from their diverse training data. However, finetuning often incurs catastrophic forgetting of skills in the pretrained model (He et al., 2021; Luo et al., 2023), which is carried over when these finetuned models are merged. As our method makes minimal change to the pretrained model, such forgetting is substantially mitigated. For instance, in the vision setting where we localize $1\%$ parameters for each task, the total parameters changed in our merged model compared with the pretrained model is only around $7\%$ for the 8 tasks as a result of minimal overlap. We evaluate the retention of pretrained capabilities with a general vision task that the pretrained CLIP model excels, namely zero-shot ImageNet classification (Deng et al., 2009), and report the results in Figure 8. Our method most effectively preserves the pretrained performance, while achieves superior performance on the eight merged tasks.
Model compression through localization.
Our localization approach enables a compact representation of the finetuned model. In our experiments, we find that localizing only $1\%$ of the total parameters recovers over $99\%$ of the performance achieved by single-task finetuning (full evaluation reported in Appendix B). The efficiency allows us to store only the masked task vectors for each task ($\gamma_{i}\odot\tau_{i}$) instead of the entire finetuned models, without a noticeable loss in performance. Given the sparsity of these masked task vectors, we can store them in Compressed Sparse Row (CSR) format (Pissanetzky, 1984; Golub & Van Loan, 2013), which drastically reduces the model size to about $1\%$ of the original. For example, a RoBERTa-base model, which typically requires $\sim 650$MB of memory to store, can be represented using only a $\sim 7$MB sparse matrix, achieving a memory reduction of $99\%$. Although we still need to store the full pretrained model, this storage overhead will be amortized with more finetuned models. This model compression, combined with the ease of update mentioned in Section 3.3, enables flexible composition of skills from multiple finetuned models with minimal storage and computational overhead.
Figure 9: Comparison of performance (left) and efficiency (right) for continual learning. Localize-and-Stitch consistently outperforms baselines with generally constant runtime for each additional task.
Continual learning.
As mentioned in Section 3.3, our approach is particularly efficient in the continual learning setting. To illustrate the efficiency, we start from merging the SST-2 and CR RoBERTa models, and incrementally add 4 more tasks to simulate a continual learning setting. The tasks are selected by representativeness (including sentiment analysis, sentence classification, NLI, etc). We compare our method against five baselines: task arithmetic, TIES, Consensus TA, Consensus TIES and continual training. Following common practice, the merging coefficients are tuned across $\{0.0,0.1,...0.9,1.0\}$. The continual training results are obtained by sequential training on each task, using the model from the previous task as the starting point. The results from Figure 9 demonstrate that: i) Performance: Localize-and-Stitch consistently outperforms the baselines, with the performance margin increasing as more tasks are involved, showing its ability to reduce task interference. ii) Runtime: The runtime for other baselines increases with each added task due to the need for hyperparameter search from scratch and performance validations, while the runtime of Localize-and-Stitch generally remains constant, reflecting its efficiency in continual learning scenarios. While continual training performs better than other model merging baselines, it still falls short of Localize-and-Stitch. This suggests a potential avenue for enhancing continual learning: instead of finetuning on top of the last task model, it may be advantageous to first finetune the pretrained model on the new task and then merge it with the existing multi-task model on the old tasks. We leave further exploration of this approach for future work.